The director does not want to change the selling price which is $80 per unit. The variable expenses are $40 per unit. Instead, he wants to increase the sales commission by $2.40 per unit. He thinks that this move with some increase in advertising would increase this year's sales by 25%. How much could the director increase this year's advertising expense and still earn the same net operating income as the previous year?

Last year's data:

Sales $1,920,000 (24000 units * $80)

Variable expenses 960,000

Contribution margin 960,000

Fixed expenses 180,000

Net operating income $780,000
Answer #1

sales Units increased = 24000*1.25 = 30,000

Selling Price increased = 80 = 80.00

Sales REvenue = 30,000*80.00= $2,400,000

increse in varible cost =30000*2.40 = 72,000

VAraible cost = 30,000*40+72000 =$1,272,000

Contibution Margin = $1,128,000

Increase in advertisemnets = Contribution - Fixed cost - Operating income = 1,128,000-180,000-780,000 = $168,000 (Answer)

Increase in Advertisements = 168,000 (Answer)